# Beyond Needle(s) in the Embodied Haystack: Environment, Architecture, and Training Considerations for Long Context Reasoning

### v2

@@ Abstract
We introduce $\infty$-Thor, a new framework for long-horizon embodied tasks that advances long-context understanding in embodied AI. $\infty$-Thor provides: (1) a generation framework for synthesizing scalable, reproducible, and unlimited long-horizon trajectories; (2) a novel embodied QA task, Needle(s) in the Embodied Haystack, where multiple scattered clues across extended trajectories test agents’ long-context reasoning ability; and (3) a long-horizon dataset and benchmark suite featuring complex tasks that span hundreds of environment steps, each paired with ground-truth action sequences. To enable this capability, we explore architectural adaptations, including interleaved Goal-State-Action modeling, context extension techniques, and Context Parallelism, to equip LLM-based agents for extreme long-context reasoning and interaction. Experimental results and analyses highlight the challenges posed by our benchmark and provide insights into training strategies and model behaviors under long-horizon conditions. Our work provides a foundation for the next generation of embodied AI systems capable of robust, long-term reasoning and planning. The datasets and code can be found at pearls-lab.github.io/infini-thor.
@@ 1 Introduction
Real-world embodied reasoning is a sequential decision-making problem requiring long-horizon planning, where task success depends on both memorizing and reasoning over multiple events that occur far apart in time. Using large pre-trained vision-language-action (VLA) models as policies for such tasks requires surpassing the key challenge of long-context reasoning. We seek to answer questions pertaining to what design choices matter in terms of environments, model architectures, and training methods when using VLA models for long-horizon embodied tasks. To this end, we develop a new framework for long-horizon tasks designed to push the boundaries of long-context understanding in embodied AI.
We introduce $\infty$-Thor, a new framework for generation, training, and evaluation of long-horizon embodied tasks. Our benchmark uniquely features tasks with a synthetic final goal, which involves multiple objects that appear at distant time steps, requiring multi-step reasoning across over hundreds of steps. Figure 1 illustrates an example: the agent observes the tomato at an early step (t=17) and the counter top much later (t=560). Then, the final task is given at t=670, which requires the agent to place the tomato on the counter top. This setup highlights the challenge of long-horizon dependency, where key objects and locations must be remembered and acted upon after hundreds of steps. Beyond these long-horizon dependencies, our framework also generates low-level robot-arm manipulation action sequences aligned with the trajectories, enabling agents to bridge from high-level reasoning to fine-grained physical execution.
@@
Figure 1: Example of the trajectory and a long-horizon embodied task generated from $\infty$-Thor. The final goal (“Put the tomato on the counter top” at t=670) requires recalling both the tomato (seen at t=17) and the counter (seen at t=560) to solved the long-horizon task. Context size refers to the input token length when converting the trajectory into the LLM input space.
@@
We further investigate various architectural considerations for embodied agents to operate under extreme sequence lengths. We show that interleaved Goal-State-Action modeling—a multimodal, goal-conditioned VLA architecture that jointly models interleaved sequences of goals, states, and actions using a LLM backbone is the most practical approach for this class of problems. Moreover, since standard LLMs are constrained by fixed context windows and cannot natively handle inputs exceeding 1M tokens, we explore long-context extension techniques such as rotary embedding scaling and positional interpolation (Chen et al., 2023; Ding et al., 2024; Peng et al., 2024). Lastly, we demonstrate how to further strengthen long-context reasoning by fine-tuning the model on extended-context inputs using Context Parallelism, a parallel training strategy that allows efficient scaling to very long sequences.
We provide comprehensive experiments and analyses, demonstrating both the challenges posed by our benchmark and the behavior of baseline models under long-horizon settings. We investigate a range of training considerations, including different configurations for fine-tuning and long-context adaptation, and evaluate their impact on model performance.
Our contributions are summarized as follows:
  • We introduce $\infty$-Thor, a new framework for generating, training, and evaluating long-horizon embodied tasks, featuring synthetic final goals that require multi-step reasoning across hundreds of steps.
  • We propose a novel embodied QA task, Needle(s) in the Embodied Haystack, requiring agents to recall and reason over multiple scattered clues across extended trajectories.
  • We release a large-scale trajectory dataset and an interactive evaluation environment to support both offline imitation learning and online policy execution in long-horizon settings.
  • We describe architectural adaptations including interleaved Goal-State-Action modeling, long-context extension and Context Parallelism, tailored for interactive embodied reasoning.
  • We present empirical results and analyses, providing insights to the current capabilities and limitations of embodied AI systems on long-horizon tasks.
@@ 2 Related Work
Embodied QA and Multimodal Needle in the Haystack Tasks. Embodied QA tasks, such as EmbodiedQA (Das et al., 2018) and MM-EGO (Ye et al., 2025), require agents to answer questions grounded in visual observations with spatial and temporal reasoning, but without active environment interaction during evaluation. Our NiEH task is also related to multimodal Needle in a Haystack (NiH) problems. While early NiH focused on textual recall in long contexts (Liu et al., 2024), recent multimodal extensions add visual inputs (Wang et al., 2024; 2025), though they remain limited to shorter contexts (up to 72K tokens) and lack embodied reasoning or temporal dependencies.
Datasets and Benchmarks for Long-horizon Embodied Tasks. Recent efforts have pushed toward long-horizon embodied tasks, where agents must complete multi-step goals with extended temporal dependencies. While ALFRED (Shridhar et al., 2020) and ALFWorld (Shridhar et al., 2021) introduced multi-step instruction-following tasks with action annotations and textual grounding, their task horizons are relatively short, typically under 50 steps. BEHAVIOR-100 (Srivastava et al., 2021) evaluates agent generalization on household activities, some of which require prolonged engagement, but mainly focus on single task. BALROG (Paglieri et al., 2024) is a benchmark for testing the agentic capabilities of long-context LLMs, but its scope is limited to games. Recently, EAI (Li et al., 2024b) proposed a generalized interface to evaluate LLMs for embodied decision making, while our framework focuses on multimodal online evaluation with real-time rewards and a new long-horizon reasoning task (NiEH).
@@ 3 $\infty$-Thor: An Environment for Generating, Training, and Evaluating Long-horizon Embodied Tasks
$\infty$-Thor features with a generation framework for synthesizing long trajectories to train and evaluate AI agents in long-horizon embodied tasks. We build $\infty$-Thor upon AI2-THOR (Kolve et al., 2017) simulator, an interactive 3D environment for embodied AI research that supports diverse scenes, objects, and agent actions. $\infty$-Thor enables the creation of unlimited trajectories with arbitrarily long, and provides an evaluation setup where agents can interact dynamically with the environment during both training and testing. This supports both offline learning by producing large-scale datasets, and online learning through direct agent-environment interaction.
Each trajectory generated by $\infty$-Thor consists of multiple task goals, such as “Put a clean sponge on a metal rack” and “Pick up the apple and place it on the microwave”, requiring grounded understanding and action to achieve the goal. At the end of each trajectory, the agent is assigned a synthetic long-horizon task that requires reasoning over entities encountered at distant time steps. For the example in Figure 1, the long-horizon task (Sub-goal #23) at step t=689, “Put the tomato on the counter top”, depends on observations made far earlier: the tomato at t=17 and the counter top at t=560.
Our generation framework can generate unlimited tasks, the trajectories can be exceptionally long, exceeding 1M context tokens or beyond when the trajectory is processed with LLMs. Moreover, $\infty$-Thor can generate low-level robot manipulation trajectories which are compatible with the ManipulaTHOR (Ehsani et al., 2021) simulator, supporting both symbolic plans and low-level controls. Successfully completing this task requires the agent to (1) memorize and integrate key environmental information over hundreds of steps, and (2) plan actions based on dependencies that are separated in time, demonstrating the need for long-context reasoning and robust spatio-temporal memory, and (3) execute low-level physical actions through extended manipulation sequences.
@@ 3.1 Static Evaluation: Needle(s) in the Embodied Haystack
Figure 2 presents examples of the two NiEH task types. In the single-evidence setting, a question is answerable based on a single observation step; in the multi-evidence setting, multiple temporally distant steps must be combined to answer the question. The NiEH testset includes diverse question types, such as binary (“yes” or “no”), “what”-, “where”-, and “how many”-style questions. These questions span a broad range of difficulty, from simple memory recall (similar to the Needle in a Haystack paradigm) to complex queries that requiring multi-step reasoning across temporally and spatially distributed evidence.
Testset Construction. We first replay the generated trajectories and collect the agent’s egocentric views, along with all objects that interact with the agent throughout the trajectories, such as objects that are picked up, moved, or even simply observed. Based on these interactions, we apply a set of rule-based templates to generate QA pairs, such as “Q. What object did you slice? A. {object name}” and “Q. Is {obeject name} on the desk? A. Yes/No”. Then, we sample questions based on the frequency to ensure diversity across object types, and annotate the GT answer steps using the replay logs.
(a) Needle in the Embodied Haystack: Single-evidence question types.
(b) Needles in the Embodied Haystack: Multi-evidence question types.
Figure 2: Example of N(s)iEH task and Ground-truth steps.
After generating QA pairs and annotating the GT steps, we cross-validate the answerability of each question with GT images using four different multimodal LLMs: LLaVA-OneVision 7B (Li et al., 2024a), Qwen2.5-VL 7B (Bai et al., 2025), Deepseek-VL 7B (Lu et al., 2024), and Pixtral 12B (Agrawal et al., 2024). Since these models are highly capable at standard visual QA, we filter out the questions that none of the four models successfully answer with GT images. At test time, the entire trajectory is treated as a Haystack and then cropped based on the GT image’s depth. Full details on templates, generation rules, and the validation scores of the four models are included in the Appendix.
Challenges in Needle(s) in the Embodied Haystack. The NiEH task introduces two key challenges for current models. First, many questions require reasoning over multiple temporally distant events. As shown in Figure 2(b), the agent moves a dish sponge from the garbage bin at $t=24$, then to the counter top, and later places it into a drawer at $t=751$. A question such as “Where was the dish sponge before you put it on the counter top?” requires the model to recall and chain together multiple actions and locations across hundreds of steps. Second, some questions demand aggregating sparse and temporally scattered evidence from long trajectories. In the second example in Figure 2(b), answering “How many times did you move the credit card?” requires the model to track and count all relevant actions occurring from the beginning to the end of the episode. These challenges highlight the need for models that can perform robust long-horizon reasoning across both time and modalities in complex embodied environments.
@@ 3.2 Constructing Long-horizon Trajectories for Interactive Evaluations
With $\infty$-Thor’s generation framework, we can synthesize long-horizon trajectories to construct training, validation, and test sets for offline learning and evaluation. Our approach builds upon a planner-based method (Kolve et al., 2017), in which we sequentially concatenate multiple single-task demonstrations into a extended trajectory, while maintaining consistency in object states and agent interactions throughout.
To generate each trajectory, we first sample a task type from one of seven predefined templates (e.g., pick two objects and place, pick and place with movable receptacle). We then sample objects that are required to perform the task, such as items to be picked up or receptacles to be interacted with. Based on the sampled task and objects, we use a classical task planner that operates on PDDL-defined domains to generate ground-truth action sequences. Only successful rollouts (re-simulated without failure) are retained, ensuring the reproducibility and reliability of our trajectories. We then concatenate these successful demonstrations to construct long-horizon sequences that span hundreds of steps.
Table 2: Dataset Statistics. “low” denotes low-level robot arm manipulation actions.
@@
# avg steps (low) 3,000+ 5,000+ 5,000+
For the final goal, the involved objects are sampled exclusively from those seen during the early 20% and the final 20% of the trajectory. This enforces a long-term temporal dependency between two objects that must be jointly referenced to complete the final task. Through this procedure, we generate 2,456/125/225 trajectories for the training, validation, and test sets, respectively.
We use a similar approach for low-level manipulation actions. For the Pick-up action, a trajectory is generated from the robot arm’s start point to the object’s target point (reversed for the Put action). Each trajectory is executed through fine-grained arm movements $(\Delta x,\Delta y,\Delta z)$, and only successful rollouts are retained. Details on task types and a pseudo-algorithm for the generation process are provided in the Appendix.
4 Architectures for Long-Horizon Vision-Language-Action Models
Figure 3: Agent–environment interaction through interleaved Goal-State-Action modeling.
Embodied agents must effectively interact with complex, dynamic environments, necessitating capabilities to interpret multimodal inputs (vision, language) and produce coherent sequences of actions. Developing such Vision-Language-Action (VLA) models is particularly challenging due to the need for seamless integration of perceptual understanding, linguistic reasoning, and action prediction. Existing VLA models either use separate encoders for vision, language, and action modules (Shridhar et al., 2020), or focus on short-horizon in constrained environments (Brohan et al., 2023; Kim et al., 2024), where decisions depend only on the most recent observation and a single instruction. While recent multimodal LLMs like LLaVA (Liu et al., 2023b), MiniGPT-4 (Zhu et al., 2023), and Llama 3.2 (AI, 2024) show strong multimodal reasoning abilities, they primarily operate on static inputs (e.g., single or few images) and lack the dynamic interactivity and memory needed for long-horizon embodied tasks involving continuous vision-language-action sequences. Moreover, many state-of-the-art models are only accessible via proprietary APIs, making them impractical for real-time, controllable embodied settings and managing long-term memory states.
We explore the potential of a multimodal LLM as a unified model for VLA modeling, utilizing an interleaved input structure of goal, state (visual observations), and action tokens, as illustrated in Figure 3. This interleaved multimodal input allows the model to process vision, language, and actions concurrently, thereby facilitating more coherent, real-time interaction modeling. Specifically, our goal-conditioned agent uses a multimodal LLM backbone trained to predict subsequent actions autoregressively, conditioned on sequences of goal and state tokens. At each timestep $t$, the environment provides a new visual observation $s_{t}$, which is encoded as state tokens and appended to the existing token stream. The model then autoregressively predicts the next action $a_{t}$ conditioned on the full history of goals, states, and previously taken actions. This action is executed in the environment (e.g., “Pick up the book”), which leads to the next observation $s_{t+1}$, continuing the perception-action loop. This interactive sequence is repeated over hundreds of steps, allowing the model to reason over temporally distant information while maintaining grounded behavior in dynamic settings. By leveraging this interleaved Goal–State–Action modeling, our architecture supports coherent decision-making across long-horizon embodied tasks.

### v3

# and Training Considerations for Long Context Reasoning

@@ Abstract
We introduce $\infty$-Thor, a new framework for long-horizon embodied tasks that advances long-context understanding in embodied AI. $\infty$-Thor provides: (1) a generation framework for synthesizing scalable, reproducible, and unlimited long-horizon trajectories; (2) a novel QA task, Needle(s) in the Embodied Haystack, where scattered clues across extended trajectories test agents’ long-context reasoning ability; and (3) a long-horizon dataset and benchmark suite featuring complex tasks that span hundreds of environment steps, each paired with ground-truth action sequences. To enable this capability, we explore architectural adaptations, including interleaved Goal-State-Action and Memory-Augmented Goal-State modeling, alongside training strategies—context extension techniques and Context Parallelism—to equip VLM-based agents for extreme long-context reasoning and interaction. Experiments highlight the challenges of our benchmark and provide insights into training strategies and model behaviors under long-horizon conditions. Furthermore, we demonstrate sim-to-real transfer: our QA dataset alone improves photorealistic benchmark performance by up to +11.2%, and we successfully integrate the framework with low-level manipulation controls. Our work provides a foundation for the next generation of embodied AI systems capable of robust, long-term reasoning and planning. The datasets and code can be found at pearls-lab.github.io/infini-thor.
Machine Learning, ICML
@@ 1 Introduction
Real-world embodied reasoning is a sequential decision-making problem requiring long-horizon planning, where task success depends on both memorizing and reasoning over multiple events that occur far apart in time. Using vision-language models (VLM) as policies for such tasks requires surpassing the key challenge of long-context reasoning. We seek to answer questions pertaining to what design choices matter in terms of environments, model architectures, and training methods when using VLMs for long-horizon embodied tasks. To this end, we develop a new framework for long-horizon tasks designed to push the boundaries of long-context understanding in embodied AI.
We introduce $\infty$-Thor, a new framework for generation, training, and evaluation of long-horizon embodied tasks. Our benchmark uniquely features tasks with a synthetic final goal, which involves multiple objects that appear at distant time steps, requiring multi-step reasoning across over hundreds of steps. As illustrated in Figure 1, an agent might observe a tomato at $t=16$ and a counter top much later at $t=560$; the final task given at $t=670$ requires the agent to retrieve the tomato and place it on the counter. This setup highlights the challenge of long-horizon dependency, where key objects and locations must be remembered and acted upon after hundreds of steps.
@@
Figure 1: Overview of $\infty$-Thor. Environment: We introduce a framework for generating long-horizon trajectories, the Needle(s) in the Embodied Haystack benchmark to test reasoning over scattered clues, and interactive online evaluation. Architecture: We investigate VLM adaptations as policy models, including Interleaved Goal-State-Action and Memory-augmented models. Training: We incorporate Context Parallelism and Context Extension for efficient long-context learning. In experiments, we demonstrate Sim2Real transfer to photorealistic benchmark tasks and integration with low-level manipulation controls.
@@
With the environment established, we investigate various architectural considerations for embodied agents to operate under extreme sequence lengths. We explore two complementary approaches: Interleaved Goal-State-Action model, which directly encodes the full perception–action history as a single multimodal sequence, and Memory-Augmented Goal-State model, which replaces the historical trajectory with text summaries or retrieved visual memories to provide long-term contextual grounding. Moreover, since standard VLMs are constrained by fixed context windows and cannot natively handle inputs exceeding 1M tokens, we evaluate long-context extension techniques, including linear interpolation, dynamic scaling, YaRN, and LongRoPE (Chen et al., 2023; Ding et al., 2024; Peng et al., 2024). Lastly, to support practical training and inference at these extended horizons, we incorporate Context Parallelism via Ring Attention (Liu et al., 2023a), enabling efficient fine-tuning on ultra-long sequences and further improving the model’s ability to reason over extended temporal contexts.
Finally, we provide comprehensive experiments and analyses, demonstrating both the challenges posed by our benchmark and the behavior of baseline models under long-horizon settings. Furthermore, we validate the real-world applicability of our framework through Sim2Real transfer, showing that our QA dataset improves performance on photorealistic benchmarks. We also demonstrate the framework’s versatility by integrating it with ManipulaTHOR (Ehsani et al., 2021) to enable low-level robot arm manipulation.
To summarize, our contributions are: (1) We introduce $\infty$-Thor, a new framework for generating, training, and evaluating long-horizon embodied tasks, featuring synthetic goals that require multi-step reasoning across hundreds of steps. (2) We propose a novel embodied QA task, Needle(s) in the Embodied Haystack, requiring agents to recall and reason over multiple scattered clues across extended trajectories. (3) We release a large-scale trajectory dataset and an interactive evaluation environment to support both offline imitation learning and online policy execution. (4) We analyze architectural adaptations (Interleaved vs. Memory-Augmented) and training strategies (Context Extension, Context Parallelism) for handling long-context inputs. (5) We present empirical results demonstrating the benefits of our framework, including successful Sim2Real transfer and integration with low-level manipulation controls.
@@ 2 Related Work
Embodied QA and Multimodal Needle in the Haystack Tasks. Embodied QA tasks, such as EmbodiedQA (Das et al., 2018) and MM-EGO (Ye et al., 2025), require agents to answer questions grounded in visual observations with spatial and temporal reasoning, but without active environment interaction during evaluation. Our NiEH task is also related to multimodal Needle in a Haystack (NiH) problems. While early NiH focused on textual recall in long contexts (Liu et al., 2024), recent multimodal extensions add visual inputs (Wang et al., 2024, 2025), though they remain limited to shorter contexts (up to 72K tokens) and lack embodied reasoning or temporal dependencies.
Datasets and Benchmarks for Long-horizon Embodied Tasks. Recent efforts have pushed toward long-horizon embodied tasks, where agents must complete multi-step goals with extended temporal dependencies. While ALFRED (Shridhar et al., 2020) and ALFWorld (Shridhar et al., 2021) introduced multi-step instruction-following tasks with action annotations and textual grounding, their task horizons are relatively short, typically under 50 steps. BEHAVIOR-100 (Srivastava et al., 2021) evaluates agent generalization on household activities, some of which require prolonged engagement, but mainly focus on single task. BALROG (Paglieri et al., 2025) is a benchmark for testing the agentic capabilities of long-context LLMs, but its scope is limited to games. Recently, EAI (Li et al., 2024b) proposed a generalized interface to evaluate LLMs for embodied decision making, while our framework focuses on multimodal online evaluation with real-time rewards and a new long-horizon reasoning task (NiEH).
@@ 3 $\infty$-Thor: An Environment for Generating, Training, and Evaluating Long-horizon Embodied Tasks
Figure 2: Example of N(s)iEH task and Ground-truth steps.
$\infty$-Thor features with a generation framework for synthesizing long trajectories to train and evaluate AI agents in long-horizon embodied tasks. We build $\infty$-Thor upon AI2-THOR (Kolve et al., 2017) simulator, an interactive 3D environment for embodied AI research that supports diverse scenes, objects, and agent actions. $\infty$-Thor enables the creation of unlimited trajectories with arbitrary length, and provides an evaluation setup where agents can interact dynamically with the environment during both training and testing. This supports both offline learning by producing large-scale datasets, and online learning through direct agent-environment interaction.
Each trajectory generated by $\infty$-Thor consists of multiple task goals, such as “Put a clean sponge on a metal rack” and “Pick up the apple and place it on the microwave”, requiring grounded understanding and action to achieve the goal. At the end of each trajectory, the agent is assigned a synthetic long-horizon task that requires reasoning over entities encountered at distant time steps. For the example in Figure 1, the long-horizon task (Sub-goal #23) at step $t=670$, “Put the tomato on the counter top”, depends on observations made far earlier: the tomato at $t=16$ and the counter top at $t=560$.
Our generation framework can generate unlimited tasks, the trajectories can be exceptionally long, exceeding 1M context tokens or beyond when the trajectory is processed with LLMs. Successfully completing this task requires the agent to (1) memorize and integrate key environmental information over hundreds of steps, and (2) plan actions based on dependencies that are separated in time, demonstrating the need for long-context reasoning and robust spatio-temporal memory.
@@ 3.1 Static Evaluation: Needle(s) in the Embodied Haystack
Figure 2 presents examples of the two NiEH task types. In the single-evidence setting, a question is answerable based on a single observation step; in the multi-evidence setting, multiple temporally distant steps must be combined to answer the question. The NiEH testset includes diverse question types, such as binary (“yes” or “no”), “what”-, “where”-, and “how many”-style questions. These questions span a broad range of difficulty, from simple memory recall (similar to the Needle in a Haystack paradigm (Liu et al., 2024)) to complex queries that requiring multi-step reasoning across temporally and spatially distributed evidence.
Testset Construction. We first replay the generated trajectories to collect egocentric views and record object interactions directly from the simulator’s metadata. Using rule-based templates, we generate QA pairs (e.g., “Q. What object did you slice? A. {object}”) and sample them based on frequency to ensure diversity. To ensure valid answerability, we cross-validate questions against ground-truth images using an ensemble of four multimodal LLMs: LLaVA-OneVision, Qwen2.5-VL, Deepseek-VL, and Pixtral (Li et al., 2024a; Bai et al., 2025; Lu et al., 2024; Agrawal et al., 2024). We filter out questions that none of these models can successfully answer. Full details on generation rules and validation are provided in Appendix A.1.
Challenges in NiEH. NiEH introduces two key challenges for current models. First, questions require reasoning over multiple temporally distant events. As shown in Figure 2(b), the agent moves a dish sponge from the garbage bin at $t=24$, then to the counter top, and later places it into a drawer at $t=751$. A question such as “Where was the dish sponge before you put it on the counter top?” requires the model to recall and chain together multiple actions and locations across hundreds of steps. Second, some questions demand aggregating sparse and temporally scattered evidence from long trajectories. In the second example in Figure 2(b), answering “How many times did you move the credit card?” requires the model to track and count all relevant actions occurring from the beginning to the end of the episode. These tasks necessitate robust long-horizon reasoning across both time and modalities in complex embodied environments.
@@ 3.2 Constructing Long-horizon Trajectories for Interactive Evaluations
With $\infty$-Thor’s generation framework, we synthesize long-horizon trajectories to construct training, validation, and test sets for offline learning and evaluation. Our approach builds upon a planner-based method (Kolve et al., 2017), in which we sequentially concatenate multiple single-task demonstrations into a extended trajectory, while maintaining consistency in object states and agent interactions throughout. To generate each trajectory, we first sample a task type from one of seven predefined templates (e.g., “pick two objects and place”, “pick and place with movable receptacle”), and objects. Based on the sampled task and objects, we use a classical task planner that operates on domains specified in the Planning Domain Definition Language (PDDL) (McDermott et al., 1998) to generate ground-truth action sequences. Only successful rollouts (re-simulated without failure) are retained, ensuring the reproducibility and reliability. We then concatenate these successful demonstrations to construct long-horizon sequences that span hundreds of steps. For the final goal, the involved objects are sampled exclusively from those seen during the early 20% and the final 20% of the trajectory. This enforces a long-term temporal dependency between two objects that must be jointly referenced to complete the final task. Details on task types and a pseudo-algorithm for the generation process are provided in Appendix A.2.
Table 2: Dataset statistics. Token lengths are calculated using the LLaVA-OneVision tokenizer. See Appendix A.3 for detailed statistics.
@@
4 Architectures and Training Considerations for Long-Horizon Embodied Reasoning
Figure 3: Overview of Policy Architectures. (a) Interleaved Goal-State-Action Model: Encodes the full history as a single continuous sequence of goal, state, and action tokens. (b) Memory-augmented Goal-State Model: Retrieves relevant past information to provide long-term context without processing the entire trajectory.
Long-horizon embodied tasks pose significant challenges due to the need to interpret multimodal inputs (vision, language) and produce coherent sequences of actions over hundreds of interactive steps. However, existing approaches either use separate encoders for vision, language, and action modules (Shridhar et al., 2020; Hafner et al., 2025), or focus on short-horizon in constrained environments (Brohan et al., 2023), where decisions depend only on the most recent observation and a single instruction. While recent VLMs (Liu et al., 2023b; Lu et al., 2024; Bai et al., 2025) show strong multimodal reasoning abilities, these models are not directly suited to long-horizon embodied tasks, which are interaction-dominant, involving hundreds of egocentric frames and continuous vision-language-action sequences. Moreover, many state-of-the-art proprietary models are only accessible via paid, rate-limited APIs, which makes large-scale, real-time experimentation in embodied settings more difficult and costly in practice (Li et al., 2024b). In this work, we explore two distinct policy architectures for leveraging a VLM as a unified embodied agent: (1) Interleaved Goal-State-Action Model and (2) Memory-Augmented Goal-State Model.
Interleaved Goal-State-Action Model treats the entire perception–action stream as a single unified sequence. At each timestep, the input consists of the task goal followed by all past visual observations and actions in an interleaved order $(g_{0},s_{0},a_{0},…,s_{t},a_{t})$. This design enables the VLM backbone to jointly reason over multimodal information and autoregressively predict the next action conditioned on the full trajectory history. By leveraging interleaved modeling, our architecture supports coherent decision-making over temporally distant information while maintaining grounded behavior in dynamic settings.
Memory-Augmented Goal-State Model incorporates compact history representations into the input context. Instead of supplying the full trajectory, the model conditions on either (1) textual summaries of past states and environment information (Figure 3-(b)-left), or (2) a retrieved set of relevant historical images selected via embedding similarity (Figure 3-(b)-right). These memory tokens provide distilled representations of long-term context, allowing the model to focus on the current observation while still retaining access to key historical information. By combining memory tokens with the current state, the model can maintain semantic grounding over long horizons without exceeding context-length constraints.